# A Calculus of Overlays

### v2

@@ 4. Embedding the $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to overlay-calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application:
@@
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches overlay-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated overlay tree.
@@
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to overlay-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is overlay-calculus.
@@
  • No function spaces. Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
  • No substitution. Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
  • No reduction rules. There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975). The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required. Section 4.1 shows that the convergence behavior of the translated overlay tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract, meaning two terms are overlay-contextually equivalent iff they have the same Böhm tree.
Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity. The overlay tree semantics achieves full abstraction with strictly simpler foundations than either.
@@ 4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the overlay tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
An overlay tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
Definition 4.1 (Overlay-convergence).
An overlay tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
@@
Theorem 4.3 (Full Abstraction).
For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx\mathcal{T}(N)$ iff for every closing context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$. Then:
$\mathcal{T}(M)\approx\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
Together, Adequacy and Full Abstraction establish that the overlay tree semantics is a computation mechanism for Böhm trees (Barendregt, 1984), the canonical fully abstract model of the lazy $\lambda$-calculus, using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
@@ 4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in overlay-calculus by overlaying new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), overlay-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3). The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: overlay definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
@@ 5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as overlay files (Yang, 2025b). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
@@
Following the Church-encoding convention (Yang, 2025b), we define a factory containing the abstract product type and its two constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
@@
Addition is a separate overlay file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
Each operation is an independent overlay that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited overlays. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
@@
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overlays}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
@@ 6.1. Emergent Phenomena
Several features of overlay-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
  1. (1)
@@
    Self-referential records with lazy evaluation. An overlay’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
  2. (2)
@@
    Open recursion with symmetric inheritance. Inheritance merges definitions from independent sources, all sharing the same inherited overlay tree. Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions. Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
  3. (3)
@@
    Commutative, idempotent merge. Inheritance in overlay-calculus is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
  4. (4)
@@
    Relational semantics. Every overlay-calculus operation is defined by recursion on overlay trees, not on individual values. When an overlay tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation. An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels. The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same. In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values. When applied to set-valued overlays $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$. This is relational semantics in action: each operation distributes over every combination of its inputs. The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism. No code change is needed to move from one to the other.
  5. (5)
@@
    Function color blindness. We implemented a web application as overlay files (Yang, 2025b), initially targeting Python’s synchronous sqlite3 and http.server modules. The business logic overlay file contained no async-specific code. Later, we replaced the FFI layer with one built on aiosqlite and starlette; the business logic overlay file did not change. This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the overlay file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution. More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.

### v4

@@ 4. Embedding the $\lambda$-Calculus
The following translation maps the $\lambda$-calculus in A-normal form (ANF) (Flanagan et al., 1993) to inheritance-calculus. In ANF, every intermediate result is bound to a name and each expression contains at most one application:
@@
Every $\lambda$-term can be mechanically converted to ANF by naming intermediate results, and every ANF term is a $\lambda$-term, so ANF is neither a restriction nor an extension of the $\lambda$-calculus. The ANF structure matches inheritance-calculus naturally: each $\mathbf{let}$-binding becomes a named record property, and each application is wrapped and projected via $\mathrm{result}.$
Let $\mathcal{T}$ denote the translation function. Each $\lambda$-abstraction introduces one scope level in the translated mixin tree.
@@
The five rules above are a complete translation from the $\lambda$-calculus (via ANF) to inheritance-calculus: every ANF term has an image, and the construction is compositional. Since the $\lambda$-calculus is Turing complete, so is inheritance-calculus.
@@
No function spaces:
Scott’s denotational semantics (Scott, 1976) constructs a reflexive domain $D$ satisfying $D\cong[D\to D]$ via an inverse limit in a category of continuous lattices. Here, the semantic domain is a product of powerset lattices, with no function spaces, no continuity conditions, and no inverse limits. The only fixed-point construction is Knaster–Tarski on powerset lattices (Appendix C), which requires only set-theoretic monotonicity, which is strictly weaker than Scott continuity.
No substitution:
Operational semantics defines $\beta$-reduction via capture-avoiding substitution, a mechanism whose subtleties motivated de Bruijn indices (de Bruijn, 1972) and explicit substitution calculi. Here, variable references are resolved by the $\mathtt{this}$ function walking the inheritance tree; no substitution occurs at any point.
No reduction rules:
There are no $\beta$-reduction steps, no evaluation contexts, and no choice between call-by-name and call-by-value (Plotkin, 1975). The semantics is observational: one queries path membership, and the mutually recursive equations unfold on demand.
The key step is that function application is encoded as record inheritance, and the resulting record is observed through path queries rather than applied as a function; no function spaces are required. Section 4.1 shows that the convergence behavior of the translated mixin tree corresponds exactly to the Böhm tree (Barendregt, 1984) of the original term: the translation is adequate, meaning a term converges iff it has a head normal form, and fully abstract with respect to $\lambda$-definable contexts, meaning two $\lambda$-terms are equivalent under the translation iff they have the same Böhm tree.
Combining these two facts, elementary machinery and full abstraction, yields a noteworthy consequence: the lazy $\lambda$-calculus admits a fully abstract semantics (w.r.t. $\lambda$-definable contexts) that lives entirely in naive set theory. The only mathematical prerequisites are powerset lattices, set-theoretic monotonicity, and the Knaster–Tarski fixed-point theorem. No continuous lattices, no inverse limits, no Scott topology, no game-theoretic machinery. Scott’s $D_{\infty}$ model (Scott, 1976) uses heavier machinery yet is only adequate, not fully abstract (Plotkin, 1975); Abramsky and McCusker’s game semantics (Abramsky and McCusker, 1995) achieves full abstraction but at the cost of even greater mathematical complexity. The mixin tree semantics achieves full abstraction with strictly simpler foundations than either.
@@ 4.1. Böhm Tree Correspondence
The translation $\mathcal{T}$ induces a semantics on the $\lambda$-calculus via the mixin tree equations (1)–(6). We show that the induced semantics corresponds precisely to the Böhm tree (Barendregt, 1984) of the original $\lambda$-term.
A mixin tree $T$ converges, written $T{\Downarrow}$, if following finitely many $\mathrm{result}$ projections from the root reaches a node with both $\mathrm{argument}$ and $\mathrm{result}$ among its properties, that is, the abstraction shape produced by $\mathcal{T}(\lambda x.\,M)$. We write $T{\Uparrow}$ if no such depth exists.
Definition 4.1 (Inheritance-convergence).
A mixin tree $T$ converges, written $T{\Downarrow}$, if there exists $n\geq 0$ such that:
@@
Theorem 4.3 (Full abstraction w.r.t. $\lambda$-definable contexts).
For closed $\lambda$-terms $M$ and $N$, define $\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)$ iff for every closing $\lambda$-calculus context $C[\cdot]$, $\mathcal{T}(C[M]){\Downarrow}\Leftrightarrow\mathcal{T}(C[N]){\Downarrow}$. Then:
$\mathcal{T}(M)\approx_{\lambda}\mathcal{T}(N)\quad\Longleftrightarrow\quad\mathrm{BT}(M)=\mathrm{BT}(N)$
Here $\approx_{\lambda}$ quantifies over $\lambda$-calculus contexts only; inheritance-calculus contexts are strictly more discriminating (Section 4.2). Together, Adequacy and Full Abstraction establish that the translation $\mathcal{T}$ is a fully faithful embedding of the lazy $\lambda$-calculus into inheritance-calculus with respect to Böhm tree equivalence (Barendregt, 1984), using only powerset lattices and the Knaster–Tarski fixed-point theorem, without the topological infrastructure of domain theory. This is stronger than Scott’s denotational semantics (Scott, 1976; Wadsworth, 1976), which is adequate but not fully abstract, as it identifies terms that observational equivalence distinguishes (Plotkin, 1975). The full proof, including the Single-Path Lemma, Substitution Lemma, and Convergence Preservation, is given in Appendix D.
The relationship between $\approx_{\lambda}$ and full inheritance-contextual equivalence is clarified by the expressive asymmetry of the next section: inheritance-calculus contexts can separate $\lambda$-terms that are Böhm-tree equivalent, so ${\approx_{\lambda}}\subsetneq{(\approx_{\mathrm{inh}}|_{\mathrm{im}(\mathcal{T})})}$. This gap is precisely the source of inheritance-calculus’s additional expressiveness (Theorem A.3).
@@ 4.2. Expressive Asymmetry
Two $\lambda$-terms that are Böhm-tree equivalent—and therefore indistinguishable in the $\lambda$-calculus—can be separated in inheritance-calculus by inheriting new observable projections onto an existing definition via open extension. By Felleisen’s expressiveness criterion (Felleisen, 1991), inheritance-calculus is strictly more expressive than the $\lambda$-calculus: the forward embedding is macro-expressible (Theorem A.2), but the reverse is not (Theorem A.3). The formal definitions, constructions, and proofs are given in Appendix A; the separating construction is itself a miniature instance of the Expression Problem, adding a new observable projection to an existing definition without modifying it.
The underlying mechanism is open extension: mixin definitions compose via open recursion, so existing definitions gain new observable properties without modification. In the $\lambda$-calculus, a function’s result is closed—there is no way to add new observable projections to an already-computed value without modifying the function’s definition.
@@ 5. Case Study: Nat Arithmetic
We implemented boolean logic and natural number arithmetic, both unary and binary, as MIXINv2 files (Yang, 2026). This section traces the unary Nat case in detail. The implementation follows a declarative object-oriented style: each concern is one file; UpperCamelCase names are constructors and operations; lowerCamelCase names are fields and parameters; and since there are no methods, all names are nouns or adjectives: $\mathrm{Plus}$ is a structure, not an action. Operations expose their result through a projection field (e.g., $\mathrm{sum}$).
@@
Following the Church-encoding convention (Yang, 2026), we define a factory containing the abstract product type and its two constructors. A $\mathrm{Zero}$ value inherits $\mathrm{Product}$; a $\mathrm{Successor}$ value inherits $\mathrm{Product}$ and exposes a $\mathrm{predecessor}$ field. The file name $\mathrm{NatData}$ is itself the top-level scope:
@@
Addition is a separate MIXINv2 file that inherits $\mathrm{NatData}$ and extends each constructor with a $\mathrm{Plus}$ operation exposing a $\mathrm{sum}$ field. The base case is $0+m=m$; the recursive case reduces $\mathrm{S}(n_{0})+m$ to $n_{0}+\mathrm{S}(m)$ by delegating to $n_{0}$’s $\mathrm{Plus}$ with an incremented addend:
@@
Each operation is an independent mixin that extends a factory’s subtree. Inheritance recursively merges subtrees that share a label, so every entity in the factory acquires all operations from all inherited mixins. Crucially, this applies to return types: $\mathrm{NatPlus}$ returns values from $\mathrm{NatFactory}$, and $\mathrm{NatEquality}$ independently adds $\mathrm{Equal}$ to the same factory. Inheriting both causes the values returned by $\mathrm{Plus}$ to carry $\mathrm{Equal}$ automatically, without modifying either file. In object algebras (d. S. Oliveira and Cook, 2012) and finally tagless interpreters (Carette et al., 2009), independently defined operations do not automatically enrich each other’s return types; additional boilerplate or type-class machinery is required.
@@
The same $\mathrm{Plus}$ and $\mathrm{Equal}$ files that operate on single values thus produce, without modification, the relational semantics of logic programming: $\mathtt{overrides}$ and $\mathtt{supers}$ collect all inheritance paths, and every operation distributes over them.
@@ 6.1. Emergent Phenomena
Several features of inheritance-calculus are primitive yet have no direct counterpart in the $\lambda$-calculus. Encoding any one of them requires pervasive, non-local changes to program structure.
Self-referential records with lazy evaluation:
@@
A mixin’s properties can reference other properties and enclosing scopes through qualified this without explicit binding. In the $\lambda$-calculus, encoding such self-referential structures requires explicit allocation of mutable or lazy references (Boudol, 2004) (e.g., Haskell-style thunks or ML-style ref cells), obscuring the declarative intent.
Open recursion with symmetric inheritance:
@@
Inheritance merges definitions from independent sources, all sharing the same inherited mixin tree. Encoding this in the $\lambda$-calculus requires maintaining an extensible dictionary of methods that can be merged from multiple directions. Cook’s denotational semantics of inheritance (Cook, 1989) devoted an entire dissertation to formalizing this pattern in a functional setting.
Commutative, idempotent merge:
@@
The inheritance operation is symmetric: a record’s elements form a set, so reordering them produces the same observable result. In the $\lambda$-calculus, function composition $(f\circ g)$ is neither commutative nor idempotent, and simulating these properties requires additional machinery (Harper and Pierce, 1991; Cardelli, 1994).
Relational semantics:
@@
Every inheritance-calculus operation is defined by recursion on mixin trees, not on individual values. When a mixin tree is viewed as a trie, the correspondence with relational algebra becomes precise: a path from the root to a leaf encodes a single value (e.g., $\mathrm{Successor}\to\mathrm{predecessor}\to\mathrm{Zero}$ encodes the natural number $1$); a trie union encodes a set of such values, i.e., a unary relation. An operation that takes two trie-union operands traverses every combination of their paths through $\mathtt{bases}$ (equation 4), which is a cross-join; the transitive closure $\mathtt{bases}^{*}$ in $\mathtt{supers}$ (equation 2) computes the union of all reachable paths; and $\mathtt{properties}$ (equation 1) projects the observable labels. The carrier is a trie rather than a flat relation, and observation is coinductive—the observer drives computation by querying progressively deeper paths—rather than Datalog’s bottom-up materialization, but the relational structure of the computation is the same. In the Case Study (Section 5), the $\mathrm{Plus}$ and $\mathrm{Equal}$ operations are defined on single Nat values. When applied to set-valued mixins $\{1,2\}$ and $\{3,4\}$, the same definitions, without any modification, produce the set of all pairwise sums $\{4,5,6\}$. This is relational semantics in action: each operation distributes over every combination of its inputs. The behavior is native: the semantic equations of Section 3 resolve every property to a set of paths, so single-valued and set-valued computation are the same mechanism. No code change is needed to move from one to the other.
Function color blindness:
@@
In a web application (Yang, 2026) initially targeting Python’s synchronous sqlite3 and http.server modules, the business logic MIXINv2 file contained no async-specific code. When the FFI layer was replaced with one built on aiosqlite and starlette, the business logic MIXINv2 file did not change. This is an instance of function color blindness (Nystrom, 2015): async and sync functions are distinct types in the host language, yet the MIXINv2 file is agnostic to the distinction because its empty property slots are filled at assembly time, with coroutine handlers under asyncio, or with direct implementations under synchronous execution. More generally, empty slots can be filled with effect handlers (Plotkin and Power, 2003; Pretnar, 2015) under a CPS interpretation, or with direct implementations under a direct-style interpretation, without explicit monad machinery.
Immunity to nonextensibility:
Programming languages typically provide a mix of extensible and nonextensible constructs. In such languages, extensibility is opt-in: frameworks must encode nonextensible constructs into extensible constructs using sophisticated mechanisms (Liang et al., 1995; Lämmel and Jones, 2003; Löh and Hinze, 2006; Swierstra, 2008; Carette et al., 2009; d. S. Oliveira and Cook, 2012; Kiselyov et al., 2013; Wu et al., 2014; Kiselyov and Ishii, 2015; Wang and d. S. Oliveira, 2016; Leijen, 2017; Poulsen and van der Rest, 2023), and code that uses nonextensible constructs cannot be extended unless rewritten in the framework. In contrast, every construct of inheritance-calculus (Section 2) is extensible via deep merge (equation 3). There is no nonextensible construct, so it is impossible to write a nonextensible program. In the Case Study (Section 5), new operations and data types are added without modifying existing files. Tree-level inheritance (Section 5) and multi-path qualified this (equation 6; Appendix B) account for this. The Expressive Asymmetry (Section 4.2; Theorem A.3) provides the formal counterpart: the mixin trees in its separating construction are mechanical translations of $\lambda$-terms via $\mathcal{T}$ (Section 4), yet a new observable projection is added to them without modification. This is the extensibility that the Expression Problem (Wadler, 1998) demands. [^3] As an untyped calculus, inheritance-calculus cannot violate the extensibility requirement of the Expression Problem but does not encounter its type-safety requirement. A typed language built on inheritance-calculus (Yang, 2026) inherits the immunity to nonextensibility and need only ensure that its type system does not introduce nonextensible constructs.